OpenStudy (anonymous):

5 5/7 + 2 47/49 A. 8 33/49 B. 8 34/49 C. 7 34/49 D. 7 33/49

OpenStudy (anonymous):

\[5*(5/7) + 2*(47/49) ?\]

OpenStudy (anonymous):

is that the question?

OpenStudy (anonymous):

ye s

OpenStudy (anonymous):

a'ight. hold on @jessikitti

OpenStudy (anonymous):

2 + 5 = 7.... 5 /7 = 35 / 49, given that you add 35/49 + 47/49 which gives you 82/49 or 1 and 33 / 49.... so the answer would be 8 and 33/ 49

OpenStudy (anonymous):

weird question but i guess @skinny23 got it ^_^